# Multiple states of two-dimensional turbulence above topography

### v1

@@ 3.3 Empirical background flow
When extracting the empirical slope $\mu_{emp}$ for each time instant, we found that its value does not vary significantly in time after an initial adjustment phase. The empirical slope $\mu_{emp}$ should depend on the energy $E$, as illustrated by SY23, and the initial wavenumber $k_{ini}$ based on our experiments. The results of $\mu_{emp}$ are shown in figure 8. We choose three thresholds for extracting $\mu_{emp}$, and find only quantitative differences. The runs with the critical energy $E_{\#}$ show that $\mu_{emp}$ is always close to zero, indicating near PV homogenization. For small-$k_{ini}$ runs, the empirical slope $\mu_{emp}$ is close to the theoretical $\mu(E)$ (horizontal dashed line) shown in figure 2$(a)$. Even at the high energy level of $2E_{\#}$, a negative $\mu_{emp}$ is consistent with the theoretical $\mu(2E_{\#})$. This is another justification of the existence of the minimum enstrophy state achievable by numerical simulations but not obtained by SY23. As the initial wavenumber increases, the magnitude of $\mu_{emp}$ tends to zero, implying a transition from a non-uniform background PV field imprinted by the low-pass-filtered topography to PV homogenization, consistent with the preceding observations from the long-term snapshots. The numerical solutions of SY23 (see vertical dotted-dashed lines) might capture a confined range of the transition regime, particularly in the low-energy cases.
@@ 4 Conclusion
In this work, we study the states of unforced, weakly-decaying two-dimensional turbulence above random rough topography, initialised by a series of monoscale fields covering a broad range of length scales and energy levels. Our observations complement those of Siegelman & Young (2023), who took a single, intermediate monoscale for initialisation into account. Some of the phenomenological descriptions of topographic turbulnece from Siegelman & Young (2023) shall be updated. The minimum-enstrophy solutions of Bretherton & Haidvogel (1976) can be reached by initial fields whose length scales are comparable with the domain size, especially when the energy is no larger than $E_{\#}$. As the initial length scale decreases, the higher initial enstrophy provides more seeds for vortex nucleation. More emerging vortices shield enstrophy from cascading toward high wavenumbers, and efficiently mix the background PV. A non-uniform background PV field imprinted by large-scale topography, at low or high energy, is thus weakened and tends to homogenization. Our results highlight the sensitivity of topographic turbulence to the length scales of initial conditions. Given that the depth-invariant mode of geostrophic turbulence responsible for driving mass and heat transports was constantly found to be energized across a range of wavenumbers (Chen, 2023), this work calls for ongoing efforts for refining parameterisations of ocean turbulence over bumpy seafloor by taking the energization length scales into account.
@@
[Acknowledgements]This work is supported by the Research Grants Council (RGC) of Hong Kong under awards Early Career Scheme 26307720 and General Research Fund 16305321, and by the Center for Ocean Research (CORE), a joint research center between Laoshan Laboratory and HKUST.

### v2

@@ 3.3 Empirical background flow
When extracting the empirical slope $\mu_{emp}$ for each time instant, we found that its value does not vary significantly in time after an initial adjustment phase. The empirical slope $\mu_{emp}$ should depend on the energy $E$, as illustrated by SY23, and the initial wavenumber $k_{ini}$ based on our experiments. The results of $\mu_{emp}$ are shown in figure 8. We choose three thresholds for extracting $\mu_{emp}$, and find only quantitative differences. The runs with the critical energy $E_{\#}$ show that $\mu_{emp}$ is always close to zero, indicating near homogenization of PV. For small-$k_{ini}$ runs, the empirical slope $\mu_{emp}$ is close to the theoretical $\mu(E)$ (horizontal dashed line) shown in figure 2$(a)$. Even at the high energy level of $2E_{\#}$, a finite and negative $\mu_{emp}$ is consistent with the theoretical $\mu(2E_{\#})$. This is another justification of the existence of the minimum enstrophy state approachable by numerical simulations but not obtained by SY23. As the initial wavenumber increases, the magnitude of $\mu_{emp}$ tends to zero, implying a transition from a non-uniform background PV field imprinted by the low-pass-filtered topography to PV homogenization, consistent with the preceding observations from the long-term snapshots. The numerical solutions of SY23 (see vertical dotted-dashed lines) might capture a confined range of the transition regime, particularly in the low-energy cases.
@@
[…]
As shown in figure 3, the initial length scale directly modulates the initial enstrophy. Thus, the dependence of the long-term states on the former can be regarded as to that on the latter. Figure 9 depicts a phase diagram of all runs in the parametric space of the energy level $E/E_{\#}$ and the initial enstrophy $Q(0)/Q_{\eta}$. Two regimes of the long-term states, namely, non-uniform and homogeneous background PV, are delineated by the empirical slope $\mu_{emp}$ in comparison against a small threshold of $0.1k_{1}^{2}$. The two regimes are roughly separated by a straight line on the low-energy branch $E<E_{\#}$. We can observe that lower-energy runs need higher initial enstrophy to approach homogeneous background PV. This can be understood by the fact that the minimum-enstrophy PV with lower energy is more non-uniform and requires stronger mixing (promoted by higher enstrophy) to approach homogenization.
@@ 4 Conclusion
In this work, we study the states of unforced, weakly-decaying two-dimensional turbulence above random rough topography, initialised by a series of monoscale fields covering a broad range of length scales and energy levels. Our observations complement those of Siegelman & Young (2023), who took a single, intermediate monoscale for initialisation into account. Some of the phenomenological descriptions of topographic turbulnece from Siegelman & Young (2023) shall be updated. The minimum-enstrophy solutions of Bretherton & Haidvogel (1976) can be approached by initial fields whose length scales are comparable with the domain size, especially when the energy is no larger than $E_{\#}$. As the initial length scale decreases, the higher initial enstrophy provides more seeds for vortex nucleation. More emerging vortices shield enstrophy from cascading toward high wavenumbers, and efficiently mix the background PV. A non-uniform background PV field imprinted by large-scale topography, at low or high energy, is thus weakened and tends to homogenization. Our results highlight the sensitivity of topographic turbulence to the length scales of initial conditions. Given that the depth-invariant mode of geostrophic turbulence responsible for driving mass and heat transports was constantly found to be energized across a range of wavenumbers (Chen, 2023), this work calls for ongoing efforts for refining parameterisations of ocean turbulence over bumpy seafloor by taking the energization length scales into account.
Our work has confined the initial conditions of topographic turbulence to monoscales. How the turbulence responds to multi-scale initial conditions, such as a power-law field, is yet to be addressed. For simplicity, if one considers a superposition of two monoscale fields with identical energy and enstrophy to those of a single monoscale field, the energy and enstrophy must be redistributed between the two scales, leading to one scale larger than and the other smaller than the single monoscale. According to our monoscale experiments, a large-scale initial field tends to the minimum-enstrophy state whose PV is typically non-uniform, whereas a small-scale counterpart tends to generate vortices that mix the background PV to homogenization. Thus, there could be a competition among different scales, in term of bringing the background PV to non-uniformity or homogenization. Which effect is dominant may depend on the assigned energy to the two monoscales, among other processes. These possibilities open a future research avenue in topographic turbulence.
@@
[Acknowledgements]This work is supported by the Research Grants Council (RGC) of Hong Kong under awards Early Career Scheme 26307720 and General Research Fund 16305321, and by the Center for Ocean Research (CORE), a joint research center between Laoshan Laboratory and HKUST. The authors are grateful to the three anonymous reviewers for their constructive comments which improved this paper a lot.